HIGH-RESOLUTION INFRARED SPECTROSCOPY OF Ge_{2}C_{3}

\underline{S. THORWIRTH}, V. LUTTER, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; T. F. GIESEN, Universität Kassel, Fachbereich 10 - Physik, 34132 Kassel, Germany; J. GAUSS, Institut für Physikalische Chemie, Universität Mainz, 55099 Mainz, Germany.

Carbon-rich systems are of great importance in diverse areas of research like material science as well as astro- and structural chemistry. Despite this relevance, our knowledge of smaller cluster units is still fragmentary, particularly with respect to investigations at high-spectral resolution in the gas phase. Unequivocal assignment of spectral features to their molecular carriers is critically dependent on predictions from high-level quantum-chemical calculations. In turn, high-resolution studies provide useful information to assess the predictive power of quantum-chemical methods. This is particularly interesting for cluster systems harboring heavy elements for which so far relatively little is known from experiment. With this contribution, we would like to present a recent gas-phase study of a polyatomic germanium-carbon cluster, linear Ge_{2}C_{3} (Ge=C=C=C=Ge), which was previously studied in an Ar matrix^{a}. The cluster was produced through laser ablation of germanium-graphite sample rods and observed in a free jet at wavelengths around 5μm. Additionally, quantum-chemical calculations of Ge_{2}C_{3} were performed at the CCSD(T) level of theory. The production and observation of Ge_{2}C_{3} suggests that many more binary clusters should be amenable to high-resolution spectroscopic techniques not only in the infrared but also in the microwave region.

aD. L. Robbins, C. M. L. Rittby, and W. R. M. Graham, J. Chem. Phys. 114, 3570 (2001).